# End-to-End Crystal Structure Prediction from Powder X-Ray Diffraction

### v2

@@ 5.2 Loss Functions
$\text{sim}({\bm{P}}_{i},{\bm{C}}_{j})=\frac{{\bm{P}}_{i}\cdot{\bm{C}}_{j}}{\|{\bm{P}}_{i}\|\|{\bm{C}}_{j}\|}$
Diffusion-Based Conditional Crystal Structure Generation Loss
Our research methodology focuses on the generation of crystal structures through a diffusion-based generative framework, leveraged by conditioning on PXRD features derived from the PXRD Feature Extractor. Utilizing the Crystal Structure Network as the backbone of generative model, this approach integrates the PXRD features to guide an iterative refinement process. The whole diffusion framework is based on DiffCSP [4], the obejective loss function is almost the same and we briefly describe it below.
For diffusion on lattice, the initial distribution can be defined as $p({\bm{L}}_{T})={\mathcal{N}}(0,{\bm{I}})$ and the diffusion forward process $q({\bm{L}}_{t}|{\bm{L}}_{t-1})={\mathcal{N}}\Big{(}{\bm{L}}_{t}|\sqrt{1-\beta_{t}}{\bm{L}}_{0},{\bm{I}}\Big{)}$ aims to add small gaussian noise into ${\bm{L}}_{t-1}$ at time $t-1$, where $\beta_{t}\in(0,1)$ control the variance. The lattice distribution at time $t$ can be derived as the probability conditional on the initial distribution:
$q({\bm{L}}_{t}|{\bm{L}}_{0})={\mathcal{N}}\Big{(}{\bm{L}}_{t}|\sqrt{\bar{\alpha}_{t}}{\bm{L}}_{0},(1-\bar{\alpha}_{t}){\bm{I}}\Big{)}$
where $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}=\prod_{s=1}^{t}(1-\beta_{s})$ can be obtained by the cosine scheduler [28].
The reverse process is given by:
@@
where ${\mathcal{M}}_{t}$ is the combination of $L_{t}$ and $F_{t}$, $\mu({\mathcal{M}}_{t})=\frac{1}{\sqrt{\alpha_{t}}}\Big{(}{\bm{L}}_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\hat{\varepsilon}_{{\bm{L}}}\Big{)},\sigma^{2}({\mathcal{M}}_{t})$$=\beta_{t}\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}$. The denoising term $\hat{\varepsilon}_{{\bm{L}}}\in\mathbb{R}^{3\times 3}$ is predicted by $f_{CSP}$. For model training, the loss function of denoising lattice is defined as:
${\mathcal{L}}_{\bm{L}}=\mathbb{E}_{{\bm{\epsilon}}_{\bm{L}}\sim{\mathcal{N}}(0,{\bm{I}}),t\sim{\mathcal{U}}(1,T)}[\|\varepsilon_{{\bm{L}}}-\hat{\varepsilon}_{{\bm{L}}}\|_{2}^{2}].$ (2)
For diffusion on fractional coordinate, the distribution at time $t$ can also be obtained by ${\bm{F}}_{t}=w({\bm{F}}_{0}+\sigma_{t}{\bm{\epsilon}}_{\bm{F}})$, where the truncation function $w(\cdot)$ is defined as $w({\bm{F}})={\bm{F}}-\lfloor{\bm{F}}\rfloor\in[0,1)^{3\times N}$, ${\bm{\epsilon}}_{\bm{F}}\in\mathbb{R}^{3\times N}$ is sampled from ${\mathcal{N}}(0,{\bm{I}})$ and $\sigma_{t}$ is the noise scale which obeys the exponential scheduler: $\sigma_{0}=0$ and $\sigma_{t}=\sigma_{1}(\frac{\sigma_{T}}{\sigma_{1}})^{\frac{t-1}{T-1}}$, if $t>0$. The whole process contains the Wrapped Normal (WN) transition [29]:
$q({\bm{F}}_{t}|{\bm{F}}_{0})\propto\sum_{{\bm{Z}}\in{\mathbb{Z}}^{3\times N}}\exp\Big{(}-\frac{\|{\bm{F}}_{t}-{\bm{F}}_{0}+{\bm{Z}}\|_{F}^{2}}{2\sigma_{t}^{2}}\Big{)}.$
The reverse process is achieved by the ancestral predictor [30, 8] with the Langevin corrector [31], while the $\hat{\varepsilon}_{{\bm{F}}}$ is generated by $f_{CSP}$. Hence, the loss function of denoising fractional coordinate is defined as:
${\mathcal{L}}_{\bm{F}}=\mathbb{E}_{{\bm{F}}_{t}\sim q({\bm{F}}_{t}|{\bm{F}}_{0}),t\sim{\mathcal{U}}(1,T)}\big{[}\lambda_{t}\|\nabla_{{\bm{F}}_{t}}\log q({\bm{F}}_{t}|{\bm{F}}_{0})-\hat{\varepsilon}_{\bm{F}}\|_{2}^{2}\big{]},$ (3)
where $\lambda_{t}=\mathbb{E}^{-1}_{{\bm{F}}_{t}}\big{[}\|\nabla_{{\bm{F}}_{t}}\log q({\bm{F}}_{t}|{\bm{F}}_{0})\|_{2}^{2}\big{]}$ is approximated via Monte-Carlo sampling, which is detailed in [4].
@@ 5.3 Implementation Details
The training of both the PXRD feature extractor and the Crystal Structure Extractor follows a rigorous protocol. The models are trained on hMOF-100 and hMOF-400, a dataset of synthesized PXRD patterns and their corresponding crystal structures. We use a batch size of 64 in hMOF-100 dataset and 32 in hMOF-400 dataset and employ the Adam optimizer with a learning rate of 1e-4 and cosine-decay scheduler. The networks are implemented in PyTorch and trained on NVIDIA V100 GPUs for 400 epochs with 40 epochs warmup for both database contrastive learning task and conditional crystal structure generation task.
The combination of these methodological components and training regimens ensures the robustness and accuracy of our approach in extracting and refining crystal structures from PXRD data. The results, as we shall demonstrate, provide compelling evidence of the effectiveness of integrating advanced neural network architectures with material science domain knowledge.
6 Data and Code Availability Statements
Upon publication, we will make both the dataset and the codebase used in this study publicly available. This action is taken to facilitate reproducibility of the results, enable further research by the academic community, and enhance the integrity of the scientific process.

### v3

@@ 4.2 Contrastive PXRD-Crystal Pretraining (CPCP) Module
$\text{sim}({\bm{P}}_{i},{\bm{C}}_{j})=\frac{{\bm{P}}_{i}\cdot{\bm{C}}_{j}}{\|{\bm{P}}_{i}\|\|{\bm{C}}_{j}\|}$ (2)
where $\text{sim}({\bm{P}}_{i},{\bm{C}}_{j})$ is the cosine similarity between PXRD and crystal structure embeddings, and $\tau$ is the temperature parameter controlling the sharpness of the distribution.
By using this approach, the model learns a shared embedding space where PXRD and crystal structure features are meaningfully aligned, thus improving its ability to interpret and predict crystal structures from PXRD data.
4.3 Conditional Crystal Structure Generation (CCSG) Module
The Conditional Crystal Structure Generation (CCSG) module employs a diffusion-based approach to generate crystal structures that are consistent with input PXRD data. The Crystal Structure Network starts with an initial noisy crystal structure and iteratively refines it to produce a final structure that aligns with the PXRD features ${\bm{P}}$. This conditioning ensures that the generated structure matches the PXRD pattern.
The forward diffusion process involves the systematic addition of Gaussian noise to the lattice parameters and fractional coordinates at each timestep, gradually increasing their randomness. Conversely, the reverse diffusion process aims to iteratively remove this noise, refining the structure towards its true form. The diffusion model is defined separately for lattice parameters ${\bm{L}}$ and fractional coordinates ${\bm{F}}$, with both processes described as follows:
Lattice Parameter Diffusion
The diffusion process for lattice parameters ${\bm{L}}$ starts with an initial Gaussian distribution:
$p({\bm{L}}_{T})=\mathcal{N}(0,\mathbf{I}),$ (3)
where ${\bm{L}}_{T}$ represents the noisy lattice state at the final diffusion step $T$. The forward diffusion process progressively adds Gaussian noise to ${\bm{L}}_{t-1}$ to obtain ${\bm{L}}_{t}$:
$q({\bm{L}}_{t}|{\bm{L}}_{t-1})=\mathcal{N}({\bm{L}}_{t}\mid\sqrt{1-\beta_{t}}{\bm{L}}_{t-1},\beta_{t}\mathbf{I}),$ (4)
where $\beta_{t}\in(0,1)$ is a variance control parameter that determines the noise level added at each step $t$. Cumulatively, the distribution of ${\bm{L}}_{t}$ conditioned on the initial lattice ${\bm{L}}_{0}$ is:
$q({\bm{L}}_{t}|{\bm{L}}_{0})=\mathcal{N}\left({\bm{L}}_{t}\mid\sqrt{\bar{\alpha}t}{\bm{L}}_{0},(1-\bar{\alpha}_{t})\mathbf{I}\right),$ (5)
where $\bar{\alpha}t=\prod{s=1}^{t}(1-\beta_{s})$, effectively capturing the aggregated effect of noise over time, often scheduled by a cosine function [47].
In practice, the input typically consists of both ${\bm{L}}$ and ${\bm{F}}$ together, rather than just ${\bm{L}}$. Therefore, the equation becomes:
@@ 4.3 Conditional Crystal Structure Generation (CCSG) Module
where ${\mathcal{M}}_{t}$ is the combination of $L_{t}$ and $F_{t}$, the mean $\mu({\bm{M}}_{t})$ and variance $\sigma^{2}({\bm{M}}_{t})$ are given by:
$\mu({\mathcal{M}}_{t})=\frac{1}{\sqrt{\alpha_{t}}}\Big{(}{\bm{L}}_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\hat{\varepsilon}_{{\bm{L}}}\Big{)},$ (6)
$\sigma^{2}({\mathcal{M}}_{t})=\beta_{t}\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}$ (7)
Here, $\hat{\varepsilon}_{{\bm{L}}}$ is the predicted noise term for lattice parameters, learned by the Crystal Structure Network.
Fractional Coordinate Diffusion
For fractional coordinates ${\bm{F}}$, the diffusion process similarly adds noise, accounting for periodic boundary conditions. The forward diffusion step for ${\bm{F}}$ at time $t$ is modeled by:
${\bm{F}}_{t}=w({\bm{F}}_{0}+\sigma_{t}{\bm{\epsilon}}_{\bm{F}}),$ (8)
where the truncation function $w(\cdot)$ ensures ${\bm{F}}$ remains within periodic boundaries, ${\bm{\epsilon}}_{\bm{F}}\sim\mathcal{N}(0,\mathbf{I})$ is the Gaussian noise added, and $\sigma_{t}$ is a noise scale parameter defined by an exponential scheduler: $\sigma_{t}=\sigma_{1}\left(\frac{\sigma_{T}}{\sigma_{1}}\right)^{\frac{t-1}{T-1}}$.
The forward distribution under this setup is given by the Wrapped Normal (WN) transition:
$q({\bm{F}}_{t}|{\bm{F}}_{0})\propto\sum_{\mathbf{Z}\in\mathbb{Z}^{3\times N}}\exp\left(-\frac{|{\bm{F}}_{t}-{\bm{F}}_{0}+\mathbf{Z}|_{F}^{2}}{2\sigma_{t}^{2}}\right).$ (9)
The reverse process aims to denoise ${\bm{F}}_{t}$ back to ${\bm{F}}_{t-1}$ using an ancestral sampling approach enhanced with Langevin dynamics, where the denoising term $\hat{\varepsilon}_{{\bm{F}}}$ is again predicted by the Crystal Structure Network.
By applying these forward and reverse processes to both lattice parameters and fractional coordinates, the model iteratively refines the noisy inputs to generate crystal structures that are consistent with the PXRD data.
4.4 Diffusion-Based Generation Loss
To optimize the CCSG module, the model minimizes the loss functions for both lattice parameters and fractional coordinates:
Lattice Loss: Minimizes the difference between predicted and true noise terms for lattice parameters:
$\mathcal{L}_{\bm{L}}=\mathbb{E}_{{\bm{\epsilon}}_{\bm{L}}\sim\mathcal{N}(0,\mathbf{I}),t\sim\mathcal{U}(1,T)}\left[|\varepsilon_{{\bm{L}}}-\hat{\varepsilon}_{{\bm{L}}}|_{2}^{2}\right].$ (10)
Fractional Coordinate Loss: Focuses on the denoising of fractional coordinates:
$\mathcal{L}_{\bm{F}}=\mathbb{E}_{{\bm{F}}_{t}\sim q({\bm{F}}_{t}|{\bm{F}}_{0}),t\sim\mathcal{U}(1,T)}\left[\lambda_{t}|\nabla_{{\bm{F}}_{t}}\log q({\bm{F}}_{t}|{\bm{F}}_{0})-\hat{\varepsilon}_{\bm{F}}|_{2}^{2}\right],$ (11)
where $\lambda_{t}=\mathbb{E}^{-1}_{{\bm{F}}_{t}}\left[|\nabla{{\bm{F}}_{t}}\log q({\bm{F}}_{t}|{\bm{F}}_{0})|_{2}^{2}\right]$ is approximated via Monte Carlo sampling [21].
The Lattice Loss $\mathcal{L}_{\bm{L}}$ minimizes the L2 norm of the difference between the predicted and actual noise components, effectively learning the noise pattern and improving denoising accuracy. The Fractional Coordinate Loss $\mathcal{L}_{\bm{F}}$ leverages gradient information to enhance the model’s sensitivity to structural variations, further refining its predictive capabilities. By optimizing these loss functions, the model learns to generate accurate crystal structures that align with PXRD data while maintaining consistency with the physical and chemical properties of the materials.
@@ 4.5 Implementation Details
XtalNet is trained on hMOF-100 and hMOF-400, datasets of synthesized PXRD patterns and their corresponding crystal structures. We employ the Adam optimizer with a learning rate of 1e-4 for CPCP module training and 1e-3 for CCSG module. We use a batch size of 64 for hMOF-100 dataset in all modules, 32 for hMOF-400 dataset in CPCP module and 16 for hMOF-400 dataset in CCSG module. The networks are implemented in PyTorch and trained on NVIDIA V100 GPUs for 400 epochs with 20 epochs warmup for all datasets and modules.
Data Availability Statement
The dataset, checkpoint and code that support the findings of this study are openly available in zenodo: https://zenodo.org/records/13629658.
Supporting Information
Supporting Information is available from the Wiley Online Library or from the author.
Acknowledgements
This work is supported by the National Natural Science Foundation of China (22125502), National Science and Technology Major Project (2022ZD0114902) and National Science Foundation of China (NSFC62276005). Conflict of Interest
The authors declare no conflict of interest.